OpenStudy (anonymous):

1/1+1/2=1/x Solve for x?

OpenStudy (zzr0ck3r):

1+1/2 = 1/x 2/2+1/2 = 1/x 3/2 = 1/x x(3/2) = 1 x = 2/3

OpenStudy (anonymous):

Thanks!